OpenStudy (anonymous):

The sum of the measures of two exterior angles of a triangle is 255. What is the measure of the third exterior angle?

OpenStudy (anonymous):

For regular polygons where n = the number of sides: Measure of one internal angle = 180°(1 - (2/n)) Sum of internal angles (for all polygons, not just regular) = 180°(n - 2) Measure of one exterior angle = 360°/n Sum of exterior angles (for all polygons, not just regular) = 360° (a) The sum of the measures of exterior angles has to add up to 360°, so if we let the measure of the third exterior angle be x we can write x + 255° = 360° x = 360° - 255° x = 105° (b) Interior angle: 180°(1 - (2/n)) = 180°(1 - (2/6)) = 120° Exterior angle: 360°/n = 360°/6 = 60°

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

No problem! :)